The Abraj Al-Bait tower has 43 fewer floors than Burj Khalifa. The two buildings together have 283 floors. Find the number of floors in each building.
Answer by ikleyn(52786)

.
B - A =  43,    (1)   (" The Abraj Al-Bait tower has 43 fewer floors than Burj Khalifa.")
B + A = 283.    (2)


---------------------Add the two equations. You will get


2B = 43 + 283 = 326  ====>  B = 326/2 = 163.


Then from eq(2),  A = 283 - B = 283 - 163 = 120.


Answer.  A"  has 120 floors;  "B"  has 163 floors.